OpenStudy (anonymous):

Write an equation of the line that passes through (8, 3) and (4, 3).

OpenStudy (anonymous):

y=3

OpenStudy (anonymous):

Find the slope first..

OpenStudy (anonymous):

It is the line parallel to x axis so, y = 3 will be the equation..

OpenStudy (anonymous):

y-y1=m(x-x1) m=y2-y1/x2-x1=0/4-8=0 y-3=0 y=3

OpenStudy (anonymous):